# Dissipative particle dynamics for coarse-grained models

### v1

@@ II.1 Limitation of the method: solute permeability
If full blocking of the solute is required, repulsive interactions between solute and solvent can be added to the model, but the resulting solvent–solute depletion interactions must taken into account, which likely requires reparametrization of solute–solute interactions.
@@ III.1 Polymer dynamics
The polymer is immersed in a DPD solvent described by the standard parameters for an aqueous solution ($\rho=3r_{\textnormal{c}}^{-3}$, $\gamma=4.5k_{\textnormal{B}}T\tau/r_{\textnormal{c}}^{2}$, $a_{ij}=78k_{\textnormal{B}}T$, $r_{\textnormal{c}}=\lambda=0.646$ nm). The coupling between the polymer beads and the DPD fluid is achieved with $\gamma_{\textnormal{s}}=5\gamma$ and $r_{\textnormal{s}}=r_{\textnormal{c}}$, which models the diffusion constant of individual monomers $D=1.1~{}\textnormal{nm}^{2}/\textnormal{ns}$ (cf. Fig. 1). The system is evolved using the velocity-Verlet integrator with time step $\Delta t=0.005\tau$.
First, we measure the diffusion constant for different polymer sizes and compare the scaling with the Zimm dynamics prediction $D\propto N^{-\nu}$, with $\nu$ the scaling exponent. At good-solvent conditions $\nu=0.588$ Rubinstein and Colby (2003). The system is a cubic box with size $L=100\lambda$ and periodic boundary conditions. Polymers of length $N$ and monomer density $\rho_{\textnormal{m}}=10^{-3}/\lambda^{3}$ are equilibrated at good solvent conditions. Bead–bead repulsion is modeled as a Lennard-Jones (WCA) interaction with size $\sigma=\lambda$, strength $\varepsilon={k_{\textnormal{B}}T}$ and cutoff $r_{\textnormal{cut}}=2^{1/6}\sigma$. We measure the mean-squared-displacement of all monomers $\langle r^{2}\rangle$ for $t=5000\tau$. The diffusion constant is measured via the slope of the average mean squared-displacement as a function of time, $6D=\partial\langle r^{2}\rangle/\partial t$; at large displacements, $\langle r^{2}\rangle>(2R_{\textnormal{g}})^{2}$.
@@
where $R_{\textnormal{g,0}}$ and $R_{\textnormal{g,col}}$ are, respectively, the initial and the final (collapsed) $R_{\textnormal{g}}$ values. We consider two initial configurations, (i) a fully expanded linear polymer ($R_{\textnormal{g},0}\sim N$) and (ii) a polymer equilibrated at good solvent conditions ($R_{\textnormal{g},0}\sim N^{\nu}$) with the scaling exponent $\nu=0.588$ Rubinstein and Colby (2003). To model the collapse, we introduce attractive interactions between monomers modeled by the Lennard-Jones potential with size $\sigma=\lambda$, strength $\varepsilon=1.25k_{\textnormal{B}}T$, and cutoff $r_{\textnormal{LJcut}}=2.5\sigma$. Simulation data shown in Fig. 6 show the scaling follows $\tau_{\textnormal{c}}\sim N^{1.42\pm 0.02}$ for an initially fully stretched polymer, while $\tau_{\textnormal{c}}\sim N^{0.95\pm 0.04}$ for an initially equilibrated polymer. This scaling is close to the analytical prediction for an expanded polymer $N^{4/3}$ and agrees with the DPD predictions for an equilibrated polymer $N^{0.98\pm 0.09}$ Guo et al. (2011).
@@
Figure 6: Polymer collapse. (a,b,c) configurations of a bead-spring polymer of size $N=200$ during collapse from an initially fully-expanded linear chain at (a) $t=500\tau$, (b) $t=2000\tau$, and (c) $t=2500\tau$. (d) DPD simulation data showing collapse timescale dependence on the polymer size $N$ for an initially expanded linear chain (circles), which follows $\sim N^{1.42\pm 0.02}$ scaling, and an initially equilibrated polymer at good solvent conditions (diamonds), which follows $\sim N^{0.95\pm 0.05}$. System size for initially expanded case, $L_{x}=N+10\lambda$, $L_{y}=L_{z}=50\lambda$, and for equilibrated polymers $L_{x}=L_{y}=L_{z}=80\lambda$. Error bars denote standard errors obtained from 5 (black circles) and 20 (blue diamonds) independent simulations.
This data on polymer diffusion and collapse timescale indicate that the proposed method faithfully reproduces the hydrodynamic coupling between a polymer and solvent.
@@ III.2 Nanochannel flow
To impose the desired boundary condition at the channel walls, we investigate a pure solvent system without ions. Implementation of solid walls within DPD simulations is not straightforward due to layering artifacts that can occur next to a flat wall Pivkin and Karniadakis (2005). A no-slip boundary condition can be imposed by introducing a layer of immobilized DPD particles at the walls Barcelos et al. (2021), however, due to repulsive interactions, the slip length dependence on the wall particle density is non-monotonic. Another possibility is to impose a drag force parallel to the wall Smiatek et al. (2008, 2009).
Here we propose an alternative strategy to impose a boundary condition by coupling the DPD fluid to the immoblized wall particles only through the thermostat with coupling strength $\gamma_{\textnormal{w}}$, which is determined by the desired slip-length. The method is very similar to using a parallel drag force Smiatek et al. (2008), but is expected to be easier to employ because it does not require a separate implantation of a wall–DPD thermostat.
@@
Figure 7: Velocity flow profile from DPD simulations (black circles) and comparison to Poiseuille profiles (red curves) given by Eq. (11). (a,b) DPD data (black) from bottom to top are obtained at $\gamma_{\textnormal{w}}/\gamma=[5,3,2,1,0.5,0.3,0.2,0.1,0.05,0.03,0.02,0.01,0.005]$ at pressure gradient $G_{x}=0.003k_{\textnormal{B}}T/\lambda$ and $w=10\lambda$. Poiseuille prediction (red curves) from bottom to top shown for slip length $L_{\textnormal{s}}/\lambda=[-0.42,-0.03,1.52,5.15,7.7,15.3,30.7,51.2]$ at viscosity $\eta=2.31k_{\textnormal{B}}T\tau/\lambda^{3}$ [Eq. (11)].
@@
with $\alpha=r_{\textnormal{c}}^{2}\gamma_{\textnormal{w}}\rho/\eta$ Smiatek et al. (2008) and the positive constants $c_{1},c_{2}$ of order unity that depend on the wall solvent interaction. We find that fitting $c_{1}=0.9$ and $c_{2}=0.13$ can be used to predict $L_{\textnormal{s}}$ to accuracy within $0.1r_{\textnormal{c}}$ (Fig. 8).
@@
Having described the channel setup and the wall interaction, we show how to simulate hydrodynamic flow of an electrolyte. Free monovalent ions are modeled as charged spheres with the short-range ion–ion repulsion modeled by the WCA potential with hydrated ion diameter $\sigma=\lambda=0.646~{}\textnormal{nm}$, and interaction strength $\varepsilon=k_{\textnormal{B}}T$. The experimentally measured value for small-ion diffusivity, $D\approx\textnormal{nm}^{2}/\textnormal{ns}$, is obtained at solvent–solute coupling strength $\gamma_{\textnormal{s}}=5\gamma$ and range $r_{\textnormal{s}}=\lambda$ (cf. Fig. 3). The electrostatic interactions are calculated using PPPM Ewald summation with real-space cutoff $r_{\textnormal{ewald}}=5\lambda$ and relative force accuracy of $10^{-4}$. Slab correction factor 3.0 is used in the $y$ coordinate to simulate fixed boundary conditions at the channel walls. Electrostatic strength is determined by the Bjerrum length $l_{\textnormal{B}}=0.71\,$nm corresponding to an aqueous solution at room temperature.
We consider an electroosmotic flow in a nanochannel of width $w=8\lambda$ under an external electric field $E_{x}$. To investigate a non-symmetric case where accurate description of convection and diffusion of ions is important, the wall contains a small charged section that covers a fraction $f_{\textnormal{w}}=0.25$ of the wall surface with charge density $\sigma_{\textnormal{q}}=0.2e_{0}/\lambda^{2}$ (see Fig. 9a). The solution contains counter-ions at density $\rho_{\textnormal{ion}}=2\sigma_{\textnormal{q}}f_{\textnormal{w}}/w$. This configuration introduces a pattern in the charge density and Poisson-Boltzmann (PB) calculations predict two distinct regimes due to the localization of counter-ions at small external fields Curk et al. (2024). This effect is pronounced at non-zero slip lengths and we use $L_{\textnormal{s}}=30\lambda$, a typical order of magnitude for a slip length of an aqueous solution Joseph and Tabeling (2005).
@@
The DPDS results are in an excellent agreement with PB simulations and clearly show two distinct flow regimes (Fig. 9). The PB calculations are performed by employing the Ludwig open-source package with electrokinetics Capuani et al. (2004) using lattice size $\Delta x=\lambda/2$, reduced viscosity $\eta^{*}=0.2$ and parameters corresponding to kinematic viscosity of water $\nu=10^{-6}~{}\textnormal{m}^{2}/\textnormal{s}$. The desired slip length is obtained via a fractional bounce-back boundary condition Wolff et al. (2012) with fraction $f_{\textnormal{bb}}$. All other parameters are the same as described for DPD. Small deviations between DPD and PB occur at the transition between the two flow regimes, which we attribute to the lack of thermal fluctuations in the PB model, as well as the lattice approximation and the associated lack of unit charge discretization in PB. However, the flow velocity in the two regimes is in perfect agreement.
@@
The method could be used with multicomponent solvents that are described by two or more different types of DPD particles Pagonabarraga and Frenkel (2001); Merabia et al. (2008). In this case, the solute–solvent coupling could be distinct among different solvents, modeling different diffusivities. In addition, the chemical potential difference of the solutes in different solvents could be introduced by soft solute–solvent interactions. While we have only considered solute models based on spherical excluded-volume interactions, such as ions or monomers in a polymers, the method can be applied to non-spherically symmetric solutes, for example, by uniformly distributing ghost particles inside a non-spherical object and coupling these ghost particles to the DPD solvent via Eqs. (2) and (3).
@@ Acknowledgements.
Implementation of the DPD-solvent in the LAMMPS open-source MD package can be achieved using the existing dpd or dpd/ext pair styles Thompson et al. (2022). Let us assume that type 1 is the solvent and type 2 is the solute. For the solute–solvent, polymer–solvent and ion–solvent interaction cases discussed in Sec. II and III (using the LJ reduced units with unit length $\lambda=r_{\textnormal{c}}=1$ and DPD particle density $\rho=3\lambda^{-3}$), the solvent properties are determined by the DPD model:
pair_coeff 1 1 dpd/ext $a_{ij}$ $\gamma$ $\gamma_{\perp}$ 1 1 1

### v2

@@ II.1 Limitation of the method: solute permeability
Another limitation of solute permeability concerns fast time-dependent changes to the flow. The timescale of the solute–solvent coupling was calculated above, $\tau_{\textnormal{BD}}=2D\tau^{2}/\lambda^{2}$, and thus any temporal momentum change on timescales $t\leq\tau_{\textnormal{BD}}$ will not be adequately transferred between solute and solvent. However, for standard DPD parameters and typical solute diffusivities we find $\tau_{\textnormal{BD}}<\tau$ (Fig. 3); the timescale of solute–solvent coupling is smaller than that of solvent–solvent coupling. Therefore, DPDS is not expected to introduce additional high-frequency limitations and the standard DPDS parameters used here are sufficient for any perturbations slower than $\tau_{\textnormal{BD}}\approx 10~{}\textnormal{ps}$, which should be adequate for most applications. If even faster response times are required, $\tau$ can be reduced by increasing the thermostat friction $\gamma$.
If permeability is not allowed, for example to model impermeable membranes, the repulsive interactions between solute and solvent can be added to the model. This is partially achieved by the standard DPD approach where solute–solvent interaction, e.g. for membrane lipids Groot and Rabone (2001) or polymers Guo, Liang, and Wang (2011), is also represented as soft repulsion [Eq. (1)]. Full blocking could be achieved by hard repulsive interaction such as Lennard-Jones or WCA. In either case, the solute–solvent conservative interactions change the equilibrium distributions of the solutes and induce entropic depletion interactions between solutes if solutes and solvent particles are not of the same size. This likely requires that the solute model is designed and parametrized for specific DPD solvent parameters.
@@ III.1 Polymer dynamics
The polymer is immersed in a DPD solvent described by the standard parameters for an aqueous solution ($\rho=3r_{\textnormal{c}}^{-3}$, $\gamma=4.5k_{\textnormal{B}}T\tau/r_{\textnormal{c}}^{2}$, $a_{ij}=78k_{\textnormal{B}}T$, $r_{\textnormal{c}}=\lambda=0.646$ nm). The coupling between the polymer beads and the DPD fluid is achieved with $\gamma_{\textnormal{s}}=5\gamma$ and $r_{\textnormal{s}}=r_{\textnormal{c}}$, which models the diffusion constant of individual monomers $D=1.1~{}\textnormal{nm}^{2}/\textnormal{ns}$ (Fig. 1). The system is evolved using the velocity-Verlet integrator with time step $\Delta t=0.005\tau$.
First, we measure the diffusion constant for different polymer sizes and compare the scaling with the Zimm dynamics prediction $D\propto N^{-\nu}$, with $\nu$ the scaling exponent; $\nu=0.588$ at good-solvent conditions Rubinstein and Colby (2003). The system is a cubic box with size $L=100\lambda$ and periodic boundary conditions. Polymers of length $N$ and monomer density $\rho_{\textnormal{m}}=10^{-3}/\lambda^{3}$ are equilibrated at good solvent conditions. Bead–bead repulsion is modeled as a Lennard-Jones (WCA) interaction with size $\sigma=\lambda$, strength $\varepsilon={k_{\textnormal{B}}T}$ and cutoff $r_{\textnormal{cut}}=2^{1/6}\sigma$. We measure the mean-squared-displacement of all monomers $\langle r^{2}\rangle$ for $t=5000\tau$. The diffusion constant is measured via the slope of the average mean squared-displacement as a function of time, $6D=\partial\langle r^{2}\rangle/\partial t$; at large displacements, $\langle r^{2}\rangle>(2R_{\textnormal{g}})^{2}$.
@@
where $R_{\textnormal{g,0}}$ and $R_{\textnormal{g,col}}$ are, respectively, the initial and the final (collapsed) $R_{\textnormal{g}}$ values. We consider two initial configurations, (i) a fully expanded linear polymer ($R_{\textnormal{g},0}\sim N$) and (ii) a polymer equilibrated at good solvent conditions ($R_{\textnormal{g},0}\sim N^{\nu}$) with the scaling exponent $\nu=0.588$. To model the collapse, we introduce attractive interactions between monomers modeled by the Lennard-Jones potential with size $\sigma=\lambda$, strength $\varepsilon=1.25k_{\textnormal{B}}T$, and cutoff $r_{\textnormal{LJcut}}=2.5\sigma$, which introduces a sudden quench in solvent quality. We can estimate the Flory-Huggins $\chi$ parameter by matching the critical point of the Lennard-Jones fluid, $T_{\textnormal{c,LJ}}\approx 1.3\varepsilon/k_{\textnormal{B}}$, to the regular solution model, yielding $\chi\approx 2T_{\textnormal{c,LJ}}/T$. For $\varepsilon=1.25k_{\textnormal{B}}T$, this results in $\chi\approx 3.25$, which is sufficiently large to drive polymer collapse.
Simulation data shown in Fig. 6 show the scaling follows $\tau_{\textnormal{c}}\sim N^{1.42\pm 0.02}$ for an initially fully stretched polymer, while $\tau_{\textnormal{c}}\sim N^{0.95\pm 0.05}$ for an initially equilibrated polymer. The first scaling exponent is close to the analytical prediction for an expanded polymer $N^{4/3}$, and the second agrees with previous DPD predictions for an equilibrated polymer Guo, Liang, and Wang (2011), $\tau_{\textnormal{c}}\sim N^{0.98\pm 0.09}$.
@@
Figure 6: Polymer collapse. (a,b,c) configurations of a bead-spring polymer of size $N=200$ during collapse from an initially fully-expanded linear chain at (a) $t=500\tau$, (b) $t=2000\tau$, and (c) $t=2500\tau$. (d) DPDS simulation data showing collapse timescale dependence on the polymer size $N$ for an initially expanded linear chain (black circles), which follows $\sim N^{1.42\pm 0.02}$ scaling, and an initially equilibrated polymer at good solvent conditions (blue diamonds), which follows $\sim N^{0.95\pm 0.05}$. For comparison, we show the collapse data using the standard DPD model with soft-repulsive beads (purple triangles). System size for initially expanded case, $L_{x}=N+10\lambda$, $L_{y}=L_{z}=50\lambda$, and for equilibrated polymers $L_{x}=L_{y}=L_{z}=80\lambda$. Error bars denote standard errors obtained from 5 (black circles) and 20 (blue diamonds and purple triangles) independent simulations.
To further validate the DPDS method we compare it with the standard DPD approach. We perform collapse simulations of a neutral polymer using the standard DPD model where the conservative monomer–monomer and monomer–solute interactions are described by the DPD potentials [Eq. (1)] and thermostat with $\gamma=4.5k_{\textnormal{B}}T\tau r_{\textnormal{c}}^{-2}$ is applied between all pairs of particles. The interaction strength is described by the prefactor $a_{ij}$, where $i$ denotes a solute and $j$ a solvent. We use $a_{jj}=a_{ij}=78k_{\textnormal{B}}T$. The polymer is equilibrated in good solvent conditions ($a_{ii}=85k_{\textnormal{B}}T$) and the collapse is then initiated by changing the monomer–monomer interaction to $a_{ii}=65k_{\textnormal{B}}T$, thereby reducing the solvent quality. The interaction difference $a_{ij}-a_{ii}=13k_{\textnormal{B}}T$ resulting in the $\chi$ parameter Groot and Warren (1997) of $\chi\approx 3.5$ that is very close to above estimate for the Lennard-Jones model. The collapse timescales using the standard DPD approach agree with the DPDS method and Lennard-Jones polymer model (Fig. 6).
This data on polymer diffusion and collapse timescale indicates that the DPDS method faithfully reproduces the hydrodynamic coupling between a polymer and a solvent. Contrary to standard DPD, the DPDS method combined with Lennard-Jones interaction (or a similar hard repulsive interaction) allows straightforward addition of point charges and simulation of polyelectrolytes Yuan and Curk (2024).
@@ III.2 Nanochannel flow
To impose the desired boundary condition at the channel walls, we investigate a pure solvent system without ions. Implementation of solid walls within DPD simulations is not straightforward due to layering artifacts that can occur next to a flat wall Pivkin and Karniadakis (2005). A no-slip boundary condition can be imposed by introducing a layer of immobilized DPD particles at the walls Barcelos et al. (2021), however, due to repulsive interactions, the slip length dependence on the wall particle density is non-monotonic. Another possibility is to impose a drag force parallel to the wall Smiatek, Allen, and Schmid (2008); Smiatek et al. (2009).
Here we propose an alternative strategy to impose a boundary condition by coupling the DPD fluid to the immoblized wall particles only through the thermostat with coupling strength $\gamma_{\textnormal{w}}$, which is determined by the desired slip-length. The method is very similar to using a parallel drag force Smiatek, Allen, and Schmid (2008), but is expected to be easier to employ because it does not require a separate implantation of a wall–DPD thermostat.
@@
Figure 7: Velocity flow profile from DPD simulations (black circles) and comparison to Poiseuille profiles (red curves) given by Eq. (11). (a,b) DPD data (black) from bottom to top are obtained at $\gamma_{\textnormal{w}}/\gamma=[5,3,2,1,0.5,0.3,0.2,0.1,0.05,0.03,0.02,0.01,0.005,0.003]$ at pressure gradient $G_{x}=0.003k_{\textnormal{B}}T/\lambda$ and $w=10\lambda$. Poiseuille prediction (red curves) from bottom to top shown for slip length $L_{\textnormal{s}}/\lambda=[-0.42,-0.03,1.52,5.15,7.7,15.3,30.7,51.2]$ at viscosity $\eta=2.31k_{\textnormal{B}}T\tau/\lambda^{3}$ [Eq. (11)].
@@
with $\alpha=r_{\textnormal{c}}^{2}\gamma_{\textnormal{w}}\rho/\eta$ Smiatek, Allen, and Schmid (2008) and the positive constants $c_{1},c_{2}$ of order unity that depend on the wall solvent interaction. We find that fitting $c_{1}=0.9$ and $c_{2}=0.13$ can be used to predict $L_{\textnormal{s}}$ to accuracy within $0.1r_{\textnormal{c}}$ (Fig. 8).
@@
Having described the channel setup and the wall interaction, we show how to simulate hydrodynamic flow of an electrolyte. Free monovalent ions are modeled as charged spheres with the short-range ion–ion repulsion modeled by the WCA potential with hydrated ion diameter $\sigma=\lambda=0.646~{}\textnormal{nm}$, and interaction strength $\varepsilon=k_{\textnormal{B}}T$. The experimentally measured value for small-ion diffusivity, $D\approx\textnormal{nm}^{2}/\textnormal{ns}$, is obtained at solvent–solute coupling strength $\gamma_{\textnormal{s}}=5\gamma$ and range $r_{\textnormal{s}}=\lambda$ (Fig. 3). The electrostatic interactions are calculated using PPPM Ewald summation with real-space cutoff $r_{\textnormal{ewald}}=5\lambda$ and relative force accuracy of $10^{-4}$. Slab correction factor 3.0 is used in the $y$ coordinate to simulate fixed boundary conditions at the channel walls. Electrostatic strength is determined by the Bjerrum length $l_{\textnormal{B}}=0.71\,$nm corresponding to an aqueous solution at room temperature.
We consider an electroosmotic flow in a nanochannel of width $w=8\lambda$ under an external electric field $E_{x}$. To investigate a non-symmetric case where accurate description of convection and diffusion of ions is important, the wall contains a small charged section that covers a fraction $f_{\textnormal{w}}=0.25$ of the wall surface with charge density $\sigma_{\textnormal{q}}=0.2e_{0}/\lambda^{2}$ (see Fig. 9a). The solution contains counter-ions at density $\rho_{\textnormal{ion}}=2\sigma_{\textnormal{q}}f_{\textnormal{w}}/w$. This configuration introduces a pattern in the charge density and Poisson-Boltzmann (PB) calculations predict two distinct regimes due to the localization of counter-ions at small external fields Curk, Leyla, and Paganobarraga (2024). This effect is pronounced at non-zero slip lengths and we use $L_{\textnormal{s}}=30\lambda$, a typical order of magnitude for a slip length of an aqueous solution Joseph and Tabeling (2005).
@@
The DPDS results are in an excellent agreement with PB simulations and clearly show two distinct flow regimes (Fig. 9). The PB calculations are performed by employing the Ludwig open-source package with electrokinetics Capuani, Pagonabarraga, and Frenkel (2004) using lattice size $\Delta x=\lambda/2$, reduced viscosity $\eta^{*}=0.2$ and parameters corresponding to kinematic viscosity of water $\nu_{\textnormal{k}}=10^{-6}~{}\textnormal{m}^{2}/\textnormal{s}$. The desired slip length is obtained via a fractional bounce-back boundary condition Wolff, Marenduzzo, and Cates (2012) with fraction $f_{\textnormal{bb}}$. All other parameters are the same as described for DPD. Small deviations between DPD and PB occur at the transition between the two flow regimes, which we attribute to the lack of thermal fluctuations in the PB model, as well as the lattice approximation and the associated lack of unit charge discretization in PB. However, the flow velocity in the two regimes is in perfect agreement.
These results demonstrate the applicability of DPDS to model hydrodynamics of electrolyte solutions. When considering highly concentrated solution, the viscosity typically increases with electrolyte concentration. This effect is captured by the DPDS method, and the rate of viscosity increase with concentration depends on the solute–solvent thermostat coupling length $r_{\textnormal{s}}$ [Eq. (6)]. Thus, the value of $r_{\textnormal{s}}$ could be chosen to obtain the desired quantitative relation between viscosity and solute concentration. Moreover, direct thermostat coupling [Eqs. (2) and (3)] between solutes could be added to further increase the viscosity of highly concentrated solutions.
@@
The method could be used with multicomponent solvents that are described by two or more different types of DPD particles Pagonabarraga and Frenkel (2001); Merabia, Bonet-Avalos, and Pagonabarraga (2008). In this case, the solute–solvent coupling could be distinct among different solvents, modeling different diffusivities. In addition, the chemical potential difference of the solutes in different solvents could be introduced by soft solute–solvent interactions. While we have only considered solute models based on spherical excluded-volume interactions, such as ions or monomers in a polymers, the method can be applied to non-spherically symmetric solutes, for example, by uniformly distributing ghost particles inside a non-spherical object and coupling these ghost particles to the DPD solvent via Eqs. (2) and (3).
@@ Acknowledgements.
Implementation of the DPDS in the LAMMPS open-source MD package can be achieved using the existing dpd or dpd/ext pair styles Thompson et al. (2022). Let us assume that type 1 is the solvent and type 2 is the solute and hybrid/overlay is used to combine DPD interactions with the solute–solute interactions. For the solute–solvent, polymer–solvent and ion–solvent interaction cases discussed in Sec. II and III (using the LJ reduced units with unit length $\lambda=r_{\textnormal{c}}=1$ and DPD particle density $\rho=3\lambda^{-3}$), the solvent properties are determined by the DPD model:
pair_coeff 1 1 dpd/ext $a_{ij}$ $\gamma$ $\gamma_{\perp}$ 1 1 1.0